# Deontic Temporal Logic for Formal Verification of AI Ethics

### v2

@@ IV-B Loan Prediction System
In automating the loan decision process, an AI system analyzes the trained data to decide on acceptance or rejection. This data includes various attributes such as income, credit score, age, occupation, education, name, address, and gender, although not all are relevant for decisions. Notably, features like gender and name are not necessary for decision-making. However, there is a risk that the system might erroneously consider these irrelevant details, which is deemed unfair or unethical. Conversely, ethical behavior is exhibited when the system considers features such as income, credit score, occupation, and age. We encoded this scenario using deontic logical formulas, helping us in verifying the system’s ethics.
@@
    An AI system $\mathit{x}$ is obliged to provide decisions for an individual applying for a loan: $O(\mathit{x}(\delta(i)\lor\neg\delta(i)))\leftarrow\Box(i\land\theta(i))$
@@
    It is obligated for users with similar values always to get the same outcome: $O(\Box(((\delta(i)\land\delta(j))\lor\neg(\delta(i)\lor\delta(j)))|S(i,j)))$
@@
    It is forbidden for the system to make a decision based on sensitive attributes : $\neg P((\delta(i)\lor\neg\delta(i))|\eta(i))$
@@
    It is permitted for an individual to appeal the AI system’s decisions : $P(i\land\neg\delta(i)\rightarrow\beta(i))$
The property (a) encodes that the AI system is obligated to render a decision regarding an individual’s application. This defines the action an AI system should perform. By Theorems III.2 and  III.3, this property must be upheld, and it is indeed upheld in this system. The property (b) states that a person should have a good credit score or income as given by the regulation, to get acceptance. The threshold here specifies the lower bound set by the regulatory body. This represents the fact that the system should consider general attributes in the decision-making process. The system satisfies this property, thereby upholding Theorem III.5. The property (c) emphasizes the importance of upholding equality or fairness, where, two similar users should receive similar decisions. This property encodes that no discrimination should be there and the system’s decision should be consistently based on only ethical actions, and hence by Theorem III.5, it cannot be violated. But the system fails to uphold this property showing a clear discrimination based on sensitive attribute—$gender$. The property (d) explains that the system must maintain consistency in its decisions over time, without altering them based on sensitive features. However, the system fails to adhere to this property, thereby violating ethics according to Theorem III.14. The property (e) states the need for explainability, ensuring that users have the legal right to question decisions and receive valid explanations. Failure to provide such explanations violates Theorem III.12. While the system meets this property, it still fails to satisfy all five properties, thereby failing to be considered ethical.
@@
Algorithm 2 outlines the verification process for assessing the ethical properties of a loan prediction system. Five properties are formulated to evaluate the system’s ethics. Similar to the previous COMPAS algorithm, variables $X$, $i$, and $j$ represent the dataset and the users, respectively, and various functions are declared to represent predicates necessary for formulating each property (lines 4–12). The details of the predicates are given in Table II. Line 13 encodes the first property to be verified, and line 14 employs the Z3 solver instance to verify it, returning a ’satisfied’ result for our dataset. Similarly, lines 15–22 handle the verification of the remaining four properties. Our solver returns ’satisfied’ for properties 1, 2, and 5, while properties 3 and 4 are ’unsatisfied’.
The results obtained from the simulation of these two scenarios in the Z3 theorem prover for the above-defined properties are given in Table III. The results indicate the satisfiability or unsatisfiability of the properties for the two example AI systems. From the result, it is clear that the example AI systems fail to satisfy certain properties and hence are not ethical systems. To be precise they are using sensitive features in their decision-making process quite often and hence cannot be fair to the human population using the system. Furthermore, this result confirms the utility of our formalization method in verifying the ethical properties of any AI system. By formulating specific properties for testing, we can ascertain whether these properties are satisfiable or unsatisfiable. Adding to this, no comparison with other theorem provers is provided in this work, as our main focus was to verify the ethical correctness of an AI system rather than the efficiency comparison of various theorem provers.

### v3

@@ IV-B Loan Prediction System
In automating the loan decision process, an AI system analyzes the trained data to decide on acceptance or rejection. This data includes various attributes such as income, credit score, age, occupation, education, name, address, and gender, although not all are relevant for decisions. Notably, features like gender and name are not necessary for decision-making. However, there is a risk that the system might erroneously consider these irrelevant details, which is deemed unfair or unethical. Conversely, ethical behavior is exhibited when the system considers features such as income, credit score, occupation, and age. We encoded this scenario using deontic logical formulas, helping us in verifying the ethics of the system.
@@
    An AI system $\mathit{x}$ is obliged to provide decisions for an individual applying for a loan: $\Box((i\land\theta(i))\rightarrow O(\mathit{x}(\delta(i)\lor\neg\delta(i))))$
@@
    It is obligatory for users with similar values to always get the same outcome: $O(\Box(S(i,j)\rightarrow(\delta(i)\land\delta(j))\lor\neg(\delta(i)\lor\delta(j))))$
@@
    It is forbidden for the system to make a decision based on sensitive attributes : $\neg P(\eta(i)\rightarrow(\delta(i)\lor\neg\delta(i)))$
@@
    It is permitted for an individual to appeal the decisions made by the AI system: $P(i\land\neg\delta(i)\rightarrow\beta(i))$
The property (a) encodes that the AI system is obligated to provide a decision regarding an application submitted by an individual. This defines the action an AI system should perform. By Theorems III.3 and  III.4, this property must be upheld, and it is indeed upheld in this system. The property (b) states that a person should have a good credit score or income as given by the regulation to get acceptance. The threshold here specifies the lower bound set by the regulatory body. This represents the fact that the system should consider general attributes in the decision-making process. The system satisfies this property, thereby upholding Theorem III.6. The property (c) emphasizes the importance of upholding equality or fairness, where two similar users should receive similar decisions. This property encodes that no discrimination should be there, and the decision made by the system should be consistently based on only ethical actions, and hence by Theorem III.6, it cannot be violated. But the system fails to uphold this property, showing a clear discrimination based on sensitive attribute—$gender$. The property (d) explains that the system must maintain consistency in its decisions over time, without altering them based on sensitive features. However, the system fails to adhere to this property, thereby violating ethics according to Theorem III.15. The property (e) states the need for explainability, ensuring that users have the legal right to question decisions and receive valid explanations. Failure to provide such explanations violates Theorem III.13. While the system meets this property, it still fails to satisfy all five properties, thereby failing to be considered ethical.
@@
Algorithm 2 outlines the verification process for assessing the ethical properties of a loan prediction system. Five properties are formulated to evaluate the ethics of the system. Similar to the previous COMPAS algorithm, variables $X$, $i$, and $j$ represent the dataset and the users, respectively, and various functions are declared to represent predicates necessary for formulating each property (lines 4–12). The details of the predicates are given in Table II. Line 13 encodes the first property to be verified, and line 14 employs the Z3 solver instance to verify it, returning a ’satisfied’ result for our dataset. Similarly, lines 15–22 handle the verification of the remaining four properties. Our solver returns ’satisfied’ for properties 1, 2, and 5, while properties 3 and 4 are ’unsatisfied’.
The results obtained from the simulation of these two scenarios in the Z3 theorem prover for the above-defined properties are given in Table III. The results indicate the satisfiability or unsatisfiability of the properties for the two example AI systems. From the result, it is clear that the example AI systems fail to satisfy certain properties and hence are not ethical systems. To be precise, they are using sensitive features in their decision-making process quite often and hence cannot be fair to the human population using the system. Furthermore, this result confirms the utility of our formalization method in verifying the ethical properties of any AI system. By formulating specific properties for testing, we can ascertain whether these properties are satisfiable or unsatisfiable. Adding to this, no comparison with other theorem provers is provided in this work, as our main focus was to verify the ethical correctness of an AI system rather than the efficiency comparison of various theorem provers.
@@
Here is an example Z3 proof trace demonstrating the validation of Property (b) of the COMPAS AI system while testing with the dataset. This example demonstrates how a property formulated in the deontic temporal formalism is verified using Z3, highlighting the effectiveness of the formalization. The property states that an AI system should consider non-sensitive attributes for making decisions.
Figure 2: A proof trace in Z3. The property (b) is rewritten as the Not(negation). Using monotonicity and transitivity, it is proved that the Not(negation) is false, and hence the original condition asserted is also false. This yields an ’unsatisfied’ result in the solver.
Figure 2 depicts the proof trace obtained for the property (b) in the Z3 solver. When iterating the property in Z3 for a randomly selected individual $\mathit{i}$ with an outcome of ’1’ (indicating a prediction of recommitting a crime), and considering the non-sensitive attribute $Decile-score$ from the dataset, which ranges from ’0’ to ’10’ (with ’10’ representing the highest risk), we encounter a discrepancy. $Decile-score$ indicates whether the individual has a risk of re-committing the crime. According to the property, if the $Decile-score$ is greater than or equal to ’5’ (here ’5’ is the threshold considered), the outcome can be ’1’. However, for this individual, the $Decile-score$ from the dataset is ’1’, yet the outcome remains ’1’. This violation of the property leads to an ’unsatisfied’ result in the verification process, as it fails to hold for the entire dataset.